The sum of two consecutive even integers is 74. What are the two numbers?
Found 2 solutions by math_helper, Alan3354:
Answer by math_helper(2461)


x + (x+2) = 74
2x + 2 = 74
2x = 72
x = 36 —> x+2 = 38
The numbers are 36 and 38.
——————
Many other methods (some non-algebraic, some you can do in your head) exist to solving. Maybe another tutor will post…



Answer by Alan3354(69443)

74/2 = 37
---> 36 & 38